A $110,000 bond bearing interest at 7.5% payable annually is bought five years before maturity to yield 7% compounded semi-annually. If the bond is redeemable at par, what is the purchase price?

The purchase price of the bond is $___?

(Round the final answer to the nearest cent as needed. Round all intermediate values to six decimal places as needed.)
Answer #1

M = $110,000, n = 5 * 2 = 10 semi-annual periods, i = 7%/2 = 3.5% (semi-annually), C = 7.5%* $110,000/2 = $4,125 (semi-annually)

P = $34,305.996956 + $77,981.069508

P = $112,287.1